OpenStudy (anonymous):

What are the period and phase shift for f(x) = tan(2x + 𝜋)?

OpenStudy (anonymous):

Somebody help please

OpenStudy (anonymous):

period formula is 2pi/k

OpenStudy (anonymous):

I'm sorry? I was absent for this lesson I have no clue what is going on.

OpenStudy (anonymous):

where k is /2/

OpenStudy (anonymous):

meaning period = 2pi/3

OpenStudy (sidsiddhartha):

take a general case, for a function like this: \[f(x)= Atan(Bx-C) +D\\period=\frac{ \pi }{ B }\\phase.shift=\frac{ C }{ B }\]

OpenStudy (sidsiddhartha):

now just compare with the above case

OpenStudy (anonymous):

So would it be, Pi/2 and (whatever that box is)/2?

OpenStudy (sidsiddhartha):

yes period is pi/2

OpenStudy (anonymous):

Awesome. Thank you all for your help!

OpenStudy (sidsiddhartha):

yw!!

OpenStudy (anonymous):

@sidsiddhartha would you mind helping with another?

OpenStudy (anonymous):

http://gyazo.com/45061329f86b44afdd8c7b42d42e3ef6 is the problem